please help me:
theta is an standard position with T(-20/-21) on its terminal ray.Find tan theta
here are the possible answers:
a)20/21
b)21/20
c)-21/21

Answer by Alan3354(69443)

theta is an standard position with T(-20/-21) on its terminal ray.Find tan theta
---------------
tan = opp/adj = y/x
tan = -21/-20
tan = 21/20